OpenStudy (mrmudd184):

will fan and medal if helped. rly need help. part 2

OpenStudy (anonymous):

List 3 values that would make this inequality true. 4n < or equal to 12 _____,_______,______

OpenStudy (photon336):

so translation: 4 times n is less than or equal to 12

OpenStudy (anonymous):

yes that is right

OpenStudy (photon336):

so you can either solve the inequality or just plug in a number that when multiplied by 4 is less than 12.

OpenStudy (marcoreus11):

^ do what @Photon336 said

OpenStudy (anonymous):

can you do the equation or draw it plz. im getting confused

OpenStudy (marcoreus11):

so plug 1 for n? is that greater or smaller than 12?

OpenStudy (marcoreus11):

u need to use 3 different number they can be equal or less but not greater

OpenStudy (photon336):

If you're stuck you can actually just solve the inequality, \[4n \le 12\] Get n by itself by dividing both sides by 4 \[\frac{ \cancel\4 }{ \cancel\4 }n \le \frac{ 12 }{ 4 }\] \[n \le 3\] @MrMudd184 what does this mean?

OpenStudy (anonymous):

3 is one answer

OpenStudy (marcoreus11):

that's one of it u need 3 answer

OpenStudy (marcoreus11):

if 3x4=12 that mean u cant use anymore number higher than 3 or else it would be bigger than 12 and make the statement false

OpenStudy (anonymous):

in fact there are infinite numbers <=3

OpenStudy (anonymous):

the next one will take me a while to key

OpenStudy (anonymous):

A gym membership requires a one-time fee of $30, and then $25 a month. How much would you spend on a hym membership for 6 months?

OpenStudy (anonymous):

30+6*25=?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

im so dumb. i thought he was telling me the equation. XD

OpenStudy (anonymous):

and i was guesing it was right

OpenStudy (anonymous):

the answer i got was 540

OpenStudy (marcoreus11):

try medaling @surjithayer or whoever u feel like helped u the best

OpenStudy (anonymous):

u help a lot, @surjithayer helped, and others. i caon only give 1 medal... :C

OpenStudy (anonymous):

can*

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

the next 1 is finding the volume

OpenStudy (anonymous):

what is the volume of 4 in., 9 1/2 in., 9 1/2 in.?

OpenStudy (marcoreus11):

i believe this is a rectangle just multiply them all

OpenStudy (anonymous):

it is a rectangle

OpenStudy (marcoreus11):

yup the formula is l x h x w so just multiply them all

OpenStudy (anonymous):

i got 361

OpenStudy (marcoreus11):

try starting a new question for a problem so it doesn't get to stuff up in here:) and yup its 361